OpenStudy (alyssajobug):

Classify this system of linear equasions 2x-6y=18 -9+x+3y A. Consistant, Dependant B. Consistant, Independant C. Inconsistant

OpenStudy (solomonzelman):

retype it.

OpenStudy (mathstudent55):

There is no equal sign in the second equation. Where should it be?

OpenStudy (alyssajobug):

\[2x -6y =18\] \[-9+x +3y\]

OpenStudy (mathstudent55):

Once again, the second line contains an expression, not an equation. You need an equal sign in the second line.

OpenStudy (alyssajobug):

Should be \[-3y =9+3y\] \[2x -6y =18\]

OpenStudy (alyssajobug):

or \[-6y =-2x +18\]

OpenStudy (mathstudent55):

The first equation is supposed to have y in both terms. There is no x in it?

OpenStudy (alyssajobug):

\[y =\frac{ 1 }{ 3 }x +18\] \[y =\frac{ x }{ 3 }+3\]

OpenStudy (alyssajobug):

Thats the solution I got but I'm not sure how to classify solutions

OpenStudy (mathstudent55):

These two equations have the same slope and different y-intercepts. They represent parallel lines. Parallel lines do not intersect. That means there is no point that is a solution to both equations. Now let's look at what your choices mean.

OpenStudy (mathstudent55):

(Notice the spelling of consistEnt, inconsistEnt, independEnt, dependEnt.) This is what these terms mean: A. Consistent, Dependent - both equations are the same equation. There is an infinite number of solutions. B. Consistent, Independent - there are two different equations whose lines intersect at one point. There is one solution. C. Inconsistent - the equations represent parallel lines. The lines do not intersect, and there is no solution.

OpenStudy (alyssajobug):

SO C. then? because they are parallel?